# SubIQ: Inverse Soft-Q Learning for Offline Imitation with Suboptimal Demonstrations

### v1

@@ 4 Practical Algorithm – SubIQ
[…]
We can adjust the scale parameter $\beta$ to regulate the conservativeness of the objective. Intuitively, if we optimize $Q$ over a lower-bounded Q-space, increasing the scale parameter $\beta$ will force each $Q(s,a)$ towards its lower bound. Consequently, when $\beta$ is sufficiently large, $\widehat{Q}^{C}$ will point-wise lower-bound $\widehat{Q}$, i.e., $\widehat{Q}^{C}(s,a)\leq\widehat{Q}(s,a)$ for all $(s,a)$.
Algorithm 1 SubIQ: Inverse soft-Q Learning with Sub-optimal Demonstrations
0:  $({\mathcal{D}}^{1},{\mathcal{D}}^{2},...,{\mathcal{D}}^{N}),(w_{1},w_{2},...,w_{N}),\overline{r}_{\eta},Q_{\psi},\pi_{\theta},\phi$
  # estimate reward reference function
  for iteration $i...N_{e}$ do
     $d\leftarrow(d^{1},d^{1},...,d^{N})\sim({\mathcal{D}}^{1},{\mathcal{D}}^{2},...,{\mathcal{D}}^{N})$
     from dataset $d$, calculate ${\mathcal{L}}(\overline{r}_{\eta})$ by (12)
     $\eta\leftarrow\eta-\nabla_{\eta}{\mathcal{L}}(\overline{r}_{\eta})$
  end for
  # train Sub-IQ
  for iteration $i...N$  do
     $d\leftarrow(d^{1},d^{1},...,d^{N})\sim({\mathcal{D}}^{1},{\mathcal{D}}^{2},...,{\mathcal{D}}^{N})$
     # Update Q function
     from dataset $d$, calculate ^$\widehat{{\mathcal{H}}}^{C}(Q_{\psi},\pi_{\theta})$ by (13)
     $\psi\leftarrow\psi+\nabla_{\psi}\widehat{{\mathcal{H}}}^{C}(Q_{\psi},\pi_{\theta})$
     # Update policy for actor-critic
     $\theta\leftarrow\theta+\nabla\left[\mathbb{E}\begin{subarray}{c}s\sim d\\ a\sim\pi_{\theta}(a|s)\end{subarray}[Q_{\psi}(s,a)-\ln(\pi_{\theta}(a|s))]\right]$
  end for
@@ 5 Experiments
Table 1: Comparison results for Mujoco and Panda-gym tasks. Values falling within the range of the best score ± 2% are emphasized in bold.
Cheetah Ant Walker Hopper Humanoid Reach Push PnP Slide
BC-E -3.2±$\pm 0.9$ 6.4$\pm$19.1 0.3$\pm$0.4 8.3$\pm$4.1 1.3$\pm$0.2 13.9$\pm$5.9 8.2$\pm$3.8 3.7$\pm$2.7 0.0$\pm$0.0
BC-O 14.2$\pm$2.9 35.2$\pm$20.1 66.1$\pm$10.8 16.7$\pm$4.0 10.6$\pm$6.3 16.2$\pm$5.5 8.8$\pm$4.5 3.9$\pm$2.7 0.1$\pm$0.3
BC-both 13.2$\pm$3.6 47.0$\pm$5.9 63.9$\pm$7.3 22.6$\pm$12.8 9.0$\pm$3.5 16.3$\pm$5.0 9.0$\pm$4.3 4.4$\pm$3.0 0.1$\pm$0.4
W-BC 12.9$\pm$2.8 47.3$\pm$6.4 58.6$\pm$9.1 20.9$\pm$6.6 19.6$\pm$19.0 15.7$\pm$5.1 8.8$\pm$4.3 3.7$\pm$2.8 0.0$\pm$0.0
TRAIL -4.1$\pm$0.3 -4.7$\pm$1.9 0.2$\pm$0.3 1.1$\pm$0.5 2.6$\pm$0.6 18.3$\pm$5.1 11.7$\pm$4.0 7.8$\pm$3.7 1.7$\pm$1.8
IQ -3.4$\pm$0.6 -3.4$\pm$1.3 0.1$\pm$0.7 0.3$\pm$0.2 2.4$\pm$0.6 97.7$\pm$2.4 26.3$\pm$10.9 18.1$\pm$12.5 0.1$\pm$0.4
SQIL -5.0$\pm$0.7 -33.8$\pm$7.4 0.2$\pm$0.2 0.2$\pm$0.0 0.9$\pm$0.1 22.1$\pm$15.1 9.6$\pm$3.3 3.2$\pm$2.9 0.1$\pm$0.3
DemoDICE 0.4$\pm$2.0 31.7$\pm$8.9 7.2$\pm$3.1 18.7$\pm$7.5 2.6$\pm$0.8 14.0$\pm$5.3 8.1$\pm$3.7 4.3$\pm$2.4 0.1$\pm$0.5
DWBC -0.2$\pm$2.5 10.4$\pm$5.0 25.1$\pm$16.3 66.0$\pm$20.9 3.7$\pm$0.3 93.4$\pm$4.3 36.9$\pm$7.4 25.0$\pm$6.3 11.6$\pm$4.4
SQRF (ours) 46.8$\pm$14.0 78.8$\pm$6.2 90.2$\pm$6.8 59.9$\pm$21.5 5.7$\pm$1.1 99.2$\pm$0.8 72.7$\pm$5.2 66.5$\pm$7.6 36.6$\pm$7.0
A-IQ (ours) 7.8$\pm$2.7 76.6$\pm$5.2 71.3$\pm$24.6 25.2$\pm$18.5 4.7$\pm$1.1 99.0$\pm$1.1 76.8$\pm$4.4 68.9$\pm$6.3 39.1$\pm$6.2
Sub-IQ (ours) 68.1$\pm$3.0 77.5$\pm$4.9 93.8$\pm$3.1 65.2$\pm$29.2 68.1$\pm$10.9 99.8$\pm$0.9 70.0$\pm$6.8 70.8$\pm$6.7 42.8$\pm$6.6
@@ 5.1 Experiment Setup
We compare our sub-IQ with state-of-the-art algorithms in offline IL with sub-optimal demonstrations: TRAIL (Yang et al., 2021), DemoDICE (Kim et al., 2021), and DWBC (Xu et al., 2022). Moreover, we also compare with other IL baselines: BC with only sub-optimal datasets (BC-O), BC with only expert data (BC-E), BC with all datasets (BC-both), and BC with fixed weight for each dataset (W-BC), SQIL (Reddy et al., 2019), and IQ-learn (Garg et al., 2021) with only expert data [^1] The performances of SQIL and IQlearn with both expert and sub-optimal data are much worse. In particular, since TRAIL, DemoDICE, and DWBC were designed to work with only two datasets (one expert and one supplementary), we combine all the sub-optimal datasets into one single supplementary dataset. Meanwhile, BC-E, BC-O, and BC-both combine all available data into a large dataset for learning, while W-BC optimizes $\sum_{i}\mathbb{E}\begin{subarray}{c}s,a\sim{\mathcal{D}}^{i}\end{subarray}\!\!-\!\!w_{i}\ln(\pi(a|s))$ We also provide comparison results with two variants of our sub-IQ: (i) A-IQ (adaptive IQ-learn), derived by retaining the first and second terms of our primary objective  (7), and (ii) SQRF (soft-Q learning with reward reference), obtained by retaining only the last term of  (7). The aim is to assess how the distribution matching and reward regularization terms help each other to achieve better performance in our sub-IQ. Here we note that the conservative Q-learning term is employed in both sub-IQ and the two variants to enhance stability. Without this term, these algorithms exhibit much worse performance and are highly unstable.
Environments and Data Generation
: We test on five Mujoco tasks  (Todorov et al., 2012) and four arm manipulation tasks from Panda-gym (Gallouédec et al., 2021). The maximum number of $(s,a)$ per trajectory is 1000 for the Mujoco and is 50 for Panda-gym tasks (descriptions in Appendix B.1). In our main experiments reported below, we test with three expertise levels (i.e., one expert and two sup-optimal datasets). Experiments with more and less expertise levels will be provided in the Appendix. The sub-optimal demonstrations have been generated by randomly adding noise to expert actions and interacting with the environments. More details of these generated datasets can be found in the Appendix B.3.
@@
The return has been normalized by the formula $\texttt{score}=\frac{\texttt{return}-\texttt{R.return}}{\texttt{E.return}-\texttt{R.return}}\times 100$ where R.return is the mean return of random policy and E.return is the mean return of the expert policy. The scores are calculated by taking the last ten evaluation scores of each seed, with at least five seeds per report.
Comparison settings.
In the following, we will conduct experiments to evaluate the following aspects: (i) performance comparison between our algorithms and other imitation learning baselines mentioned earlier, (ii) the influence of augmenting expert data, and (iii) the impact of sub-optimal demonstrations. Additional experiments can be found in the Appendix.
@@ 5.2 Main Comparison Results
Table 1 shows comparison results for our methods and 9 different baselines across five Mujoco tasks and four Panda-gym tasks. From lowest to highest expertise levels, we randomly sample (25k-10k-1k) trajectories for Mujoco tasks and (10k-5k-100) transitions for Panda-gym tasks for every seed (details of these three-level dataset are provided in Appendix B.3). In general, Sub-IQ performs the best on 8 out of the 9 tasks, and is competitive for the other task. Surprisingly, SQRF’s performance is quite comparable compared to Sub-IQ for several tasks. However, it faces difficulties in Humanoid tasks where its performance is remarkably low. None of the BC variants do well, except for the Walker (easiest Mujoco task). Furthermore, TRAIL’s performance is poor, indicating that it struggles to learn accurate dynamics using sub-optimal data. In addition, SQIL and IQ-learn also struggle with limited expert data. Moreover, our SQRF is better than SQIL, highlighting the benefits of learning with reward references instead of simply assigning +1 to every expert transition.
5.3 Augmented Expert Demonstrations
@@
Push
PnP
Ant
Push
PnP
We aim to assess the performance of sub-IQ with augmented expert data. To this end, we use two Mujoco and two Panda-gym tasks, keeping the same sup-optimal data and add more expert demonstrations to the training sets. The comparison results are reported in Figure 1. For the Mujoco tasks, which are more difficult, adding more expert trajectories significantly enhances the performance of all the algorithms. However, for the two Panda-gym tasks, the influence of adding more expert data appears to be less significant in Push and completely absent in PnP. This would be because expert trajectories in these tasks are typically short, consisting of only 2-7 transitions. Hence, a larger quantity of additional expert data may be required to enhance performance.
5.4 Impact of Sup-optimal Demonstrations
Cheetah
Ant
Push
PnP
Ant
Push
PnP
We present numerical results to assess the impact of sub-optimal data on the performance of our sub-IQ and other baselines. To this end, we keep the same expert dataset and adjust the non-expert data used in Table 1. The performance is reported in Fig.2. It is evident that reducing the amount of data in sub-optimal datasets can result in a significant degradation in the performance of our Sub-IQ and the other baselines. Conversely, adding more sub-optimal data can enhance overall stability and lead to improved performance. More detailed results can be found in Appendix B.4.
6 Conclusion
In this paper, we have developed sub-IQ, a novel non-adversarial inverse soft-Q learning algorithm for offline imitation learning from expert and sub-optimal demonstrations. We have demonstrated that our algorithm possesses several favorable properties, contributing to its well-behaved, stable, and scalable nature. Additionally, we have devised a preference-based loss function to automate the estimation of reward reference values. We have provided extensive experiments based on several benchmark tasks, demonstrating the ability of our sub-IQ algorithm to leverage both expert and non-expert data to achieve superior performance compared to state-of-the-art algorithms.
7 Broader Impacts
This paper investigates offline imitation learning with supplementary data, and our algorithms are tested on simulated tasks. Although this paper does not have immediate, direct social impacts, the potential practical applications of our research could lead to positive societal changes. By extending the applicability of imitation learning algorithms, our work may support the development of more efficient solutions in areas like robotics, autonomous vehicles, education, and healthcare. However, it is important to acknowledge that the misuse of such technology could result in negative consequences, such as manipulating information to influence people’s behavior or privacy. Hence, it is imperative to remain vigilant in ensuring that imitation learning techniques are utilized in a responsible and ethical manner.
@@ Appendix A Missing Proofs
where ${\mathcal{T}}^{\pi}[Q](s,a))=Q(s,a)-\gamma\mathbb{E}_{s^{\prime}\sim P(s^{\prime}|s,a)}[V^{\pi}(s^{\prime})]$ and $V^{\pi}(s)=\mathbb{E}_{a\sim\pi(a|s)}[Qs,a)-\log\pi(a|s)]$. We see that ${\mathcal{T}}^{\pi}[Q](s,a)$ is linear in $Q$ for any $(s,a)$. Thus, the first and second terms of $J(Q,\pi)$ in (14) are linear in $Q$. The last term of (14) involves a sum of squares of linear functions of $Q$, which are convex. So, $J(Q,\pi)$ is concave in $Q$.
@@
[…]
@@
[…]
@@
$=\sum_{i\in[N]}\sum_{(s,a),(s^{\prime},a^{\prime})\in{\mathcal{D}}^{i}}(\overline{r}(s,a)-\overline{r}(s^{\prime},a^{\prime}))^{2}-\sum_{\begin{subarray}{c}h,k\in[N],h<k\\ \tau_{i}\in{\mathcal{D}}^{h},\tau_{j}\in{\mathcal{D}}^{k}\end{subarray}}\Big{(}R(\tau_{j})-{\ln\left(\exp(R(\tau_{j}))+\exp(R(\tau_{i}))\right)}+\phi(\overline{r})\Big{)}$ (24)
We then see that the first term is a sum of squares of linear functions of $\overline{r}$, thus is strictly convex. Moreover, since $R(\tau_{i})$ is linear in $\overline{r}$ for any $\tau_{i}$ , the term $\ln(\exp(R(\tau_{i}))+\exp(R(\tau_{j})))$ has a log-sum-exp form. So this term is convex as well (Boyd & Vandenberghe, 2004). Putting all together we see that ${\mathcal{L}}(\overline{r})$ is stricly convex in $\overline{r}$ as desired. ∎
@@
[…]
@@
$\stackrel{{\scriptstyle(a)}}{{\geq}}-\beta\sum_{\begin{subarray}{c}s\sim{\mathcal{D}}\\ a\sim\mu(a|s)\end{subarray}}\widehat{Q}(s,a)+\widehat{{\mathcal{H}}}(\widehat{Q},\pi)$ (30)
@@
Appendix B Experimental Details
B.1 Environment
B.1.1 Mujoco
MuJoCo gym environments (Todorov et al., 2012) like HalfCheetah, Ant, Walker2d, Hopper, and Humanoid are integral to the field of reinforcement learning (RL), particularly in the domain of continuous control and robotics:

### v2

@@ 3.4.3 Conservative soft-Q learning
Let $\widehat{Q}=\text{argmax}_{Q}\widehat{{\mathcal{H}}}(Q,\pi)$ and $\widehat{Q}^{C}=\text{argmax}_{Q}\widehat{{\mathcal{H}}}^{C}(Q,\pi)$, we have $\sum_{\begin{subarray}{c}s\sim{\mathcal{D}}\\ a\sim\mu(a|s)\end{subarray}}\widehat{Q}^{C}(s,a)\leq\sum_{\begin{subarray}{c}s\sim{\mathcal{D}}\\ a\sim\mu(a|s)\end{subarray}}\widehat{Q}(s,a).$
We can adjust the scale parameter $\beta$ in Equation 8 to regulate the conservativeness of the objective. Intuitively, if we optimize $Q$ over a lower-bounded Q-space, increasing the scale parameter $\beta$ will force each $Q(s,a)$ towards its lower bound. Consequently, when $\beta$ is sufficiently large, $\widehat{Q}^{C}$ will point-wise lower-bound $\widehat{Q}$, i.e., $\widehat{Q}^{C}(s,a)\leq\widehat{Q}(s,a)$ for all $(s,a)$.
@@ 4 Experiments
Mujoco Panda-gym
Cheetah Ant Humanoid Push PnP Slide Avg
BC-E -3.2±$\pm 0.9$ 6.4$\pm$19.1 1.3$\pm$0.2 8.2$\pm$3.8 3.7$\pm$2.7 0.0$\pm$0.0 2.7
BC-O 14.2$\pm$2.9 35.2$\pm$20.1 10.6$\pm$6.3 8.8$\pm$4.5 3.9$\pm$2.7 0.1$\pm$0.3 12.1
BC-both 13.2$\pm$3.6 47.0$\pm$5.9 9.0$\pm$3.5 9.0$\pm$4.3 4.4$\pm$3.0 0.1$\pm$0.4 13.8
W-BC 12.9$\pm$2.8 47.3$\pm$6.4 19.6$\pm$19.0 8.8$\pm$4.3 3.7$\pm$2.8 0.0$\pm$0.0 15.4
TRAIL -4.1$\pm$0.3 -4.7$\pm$1.9 2.6$\pm$0.6 11.7$\pm$4.0 7.8$\pm$3.7 1.7$\pm$1.8 3.9
IQ-E -3.4$\pm$0.6 -3.4$\pm$1.3 2.4$\pm$0.6 26.3$\pm$10.9 18.1$\pm$12.5 0.1$\pm$0.4 6.7
IQ-both -6.1$\pm$1.4 -58.2$\pm$0.0 0.8$\pm$0.0 8.3$\pm$3.9 3.8$\pm$3.3 0.0$\pm$0.2 -8.6
SQIL-E -5.0$\pm$0.7 -33.8$\pm$7.4 0.9$\pm$0.1 9.6$\pm$3.3 3.2$\pm$2.9 0.1$\pm$0.3 -4.2
SQIL-both -5.6$\pm$0.5 -58.0$\pm$0.4 0.8$\pm$0.0 8.2$\pm$3.8 3.3$\pm$2.3 0.1$\pm$0.3 -12.6
DemoDICE 0.4$\pm$2.0 31.7$\pm$8.9 2.6$\pm$0.8 8.1$\pm$3.7 4.3$\pm$2.4 0.1$\pm$0.5 7.9
DWBC -0.2$\pm$2.5 10.4$\pm$5.0 3.7$\pm$0.3 36.9$\pm$7.4 25.0$\pm$6.3 11.6$\pm$4.4 14.6
SPRINQL (ours) 73.6$\pm$4.3 77.0$\pm$5.6 82.9$\pm$11.2 72.0$\pm$5.3 63.2$\pm$6.4 37.7$\pm$6.6 67.7
Table 1: Comparison results for three Mujoco and three Panda-gym tasks.
@@ 4.1 Experiment Setup
We compare our SPRINQL with SOTA algorithms in offline IL with sub-optimal demonstrations: TRAIL [38], DemoDICE [21], and DWBC [37]. Moreover, we also compare with other straightforward baselines: BC with only sub-optimal datasets (BC-O), BC with only expert data (BC-E), BC with all datasets (BC-both), and BC with fixed weight for each dataset (W-BC), SQIL [29], and IQ-learn [13] with only expert data. In particular, since TRAIL, DemoDICE, and DWBC were designed to work with only two datasets (one expert and one supplementary), in problems with multiple sub-optimal datasets, we combine all the sub-optimal datasets into one single supplementary dataset. Meanwhile, BC-E, BC-O, and BC-both combine all available data into a large dataset for learning, while W-BC optimizes $\sum_{i}\mathbb{E}\begin{subarray}{c}s,a\sim{\mathcal{D}}^{i}\end{subarray}\!\!-\!\!w_{i}\ln(\pi(a|s))$, where $w_{i}$ are our weight parameters. T-REX [5] is a PPO-based IL algorithm that can work with ranked demonstrations. It is, however, not suitable for our offline setting, so we do not include it in the main comparisons. Nevertheless, we will later conduct an ablation study to compare SPRINQL with an adapted version of T-REX. Moreover, [15] developed IPL for learning from offline preference data. This approach requires comparisons between every pair of trajectories, thus is not suitable for our context.
Environments and Data Generation:
We test on five Mujoco tasks  [32] and four arm manipulation tasks from Panda-gym [12]. The maximum number transition of $(s,a)$ per trajectory is 1000 (or 1k for short) for the Mujoco and is 50 for Panda-gym tasks (descriptions in Appendix B.3). The sub-optimal demonstrations have been generated by randomly adding noise to expert actions and interacting with the environments. We generated large datasets of expert and non-expert demonstrations. For each seed, we randomly sample subsets of demonstrations for testing. This approach allows us to test with different datasets across seeds, rather than using fixed datasets for all seeds as in previous works. More details of these generated databases can be found in the Appendix 5.
@@
The return is normalized by $\texttt{score}=\frac{\texttt{return}-\texttt{R.return}}{\texttt{E.return}-\texttt{R.return}}\times 100$ where R.return is the mean return of random policy and E.return is the mean return of the expert policy. The scores are calculated by taking the last ten evaluation scores of each seed, with five seeds per report.
Experimental Concerns.
Throughout the experiments, we aim to answer the following questions: (Q1) How does SPRINQL perform compared to other baselines? (Q2) How do the distribution matching and reward regularization terms impact the performance of SPRINQL? (Q3) What happens if we augment (or reduce) the expert data while maintaining the sub-optimal datasets? (Q4) What happens if we augment (or reduce) the sub-optimal data while maintaining the expert dataset? (Q5) How does the conservative term help in our approach? (Q6) How does increasing $N$ (the number of expertise levels) affect the performance of SPRINQL? (Q7) Does the preference-based weight learning approach provide good values for the weights $w_{i}$? (Q8) How does SPRINQL perform in recovering the ground-truth reward function?
@@ 4.2 Main Comparison Results
In this section, we provide comparison results to answer (Q1) with three datasets (i.e., $N=3$). Additional comparison results for $N=2$ can be found in the appendix. From lowest to highest expertise levels, we randomly sample (25k-10k-1k) transitions for Mujoco tasks and (10k-5k-100) transitions for Panda-gym tasks for every seed (details of these three-level dataset are provided in Appendix 5). Table 1 shows comparison results across 3 Mujoco tasks and 3 Panda-gym tasks (the full results for all the nine environments are provided in Appendix C.1). In general, SPRINQL significantly outperforms other baselines on all the tasks.
4.3 Ablation Study - No Distribution Matching and No Reward Regularizer
We aim to assess the importance of the distribution matching and reward regularizer terms in our objective (Q2). To this end, we conduct an ablation study comparing SPRINQL with two variants: (i) noReg-SPRINQL, derived by removing the reward regularizer term from  (6), and (ii) noDM-SPRINQL, obtained by removing the distribution matching term from  (6). Here, we note that the noDM-SPRINQL performs Q-learning using the reward reference function, this can viewed as an adaption of the T-REX algorithm [5] to our offline setting. The conservative Q-learning term is employed in the SPRINQL and the two variants to enhance stability. The comparisons for $N=2$ and $N=3$ on five Mujoco tasks are shown in Figure 1 (the full comparison results for all tasks are provided in the appendix). These results clearly show that SPRINQL outperforms the other variants, indicating the value of both terms in our objective function.
@@ 4.3 Ablation Study - No Distribution Matching and No Reward Regularizer
Walker
Hopper
Humanoid
Figure 1: Comparison of three variants of SPRINQL across five Mujoco environments.
4.4 Other Experiments
Experiments addressing the other questions are provided in the appendix. Specifically, Sections C.1 and C.2 provide full comparison results for all the Mujoco and Panda-gym tasks for three and two datasets (i.e., $N=3$ and $N=2$), complementing the answer to Q1. Section C.3 provides the learning curves of the three variants considered in Section 4.3 above (answering Q2). Section C.4 provides experiments to answer Q3 (what would happen if we augment the expert dataset?) and Section C.5 addresses Q4 (what would happen if we augment the sub-optimal dataset?). Section C.6 experimentally shows the impact of the conservative term in our approaches (i.e., Q5). Section C.7 reports the performance of SPRINQL with varying numbers of expertise levels $N$ (i.e., Q6). Section C.8 addresses Q7, and Section C.9 shows how SPRINQL performs in terms of reward recovering (i.e. Q8). In addition, Section C.10 reports the distributions of the reference rewards.
Concretely, our extensive experiments reveal the following: (i) SPRINQL outperforms other baselines with two, three, or even larger numbers of datasets; (ii) the conservative term, distribution matching, and reward regularizer terms are essential to our objective—all three significantly contribute to the success of SPRINQL; (iii) the preference-based weight learning provides good estimates for the weights $w_{i}$; and (iv) SPRINQL performs well in recovering rewards, showing a high positive correlation with the ground-truth rewards, justifying the use of our method for IRL.
5 Conclusion and Limitations
(Conclusion) We have developed SPRINQL, a novel non-adversarial inverse soft-Q learning algorithm for offline imitation learning from expert and sub-optimal demonstrations. We have demonstrated that our algorithm possesses several favorable properties, contributing to its well-behaved, stable, and scalable nature. Additionally, we have devised a preference-based loss function to automate the estimation of reward reference values. We have provided extensive experiments based on several benchmark tasks, demonstrating the ability of our SPRINQL algorithm to leverage both expert and non-expert data to achieve superior performance compared to state-of-the-art algorithms. (Limitations) Some limitations of this work include: (i) SPRINQL (and other baselines) still requires a large amount of sub-optimal datasets with well-identified expertise levels to learn effectively, (ii) there is a lack of theoretical investigation on how the sizes of the expert and non-expert datasets affect the performance of Q-learning, which we find challenging to address, and (iii) it lacks a theoretical exploration of how the reward regularizer term enhances the distribution matching term when expert samples are low—this question is relevant and interesting but also challenging to address. These limitations will pave the way for our future work.
@@ Appendix A Missing Proofs
We provide proofs for the theoritical results claimed in the main paper.
@@
where ${\mathcal{T}}^{\pi}[Q](s,a))=Q(s,a)-\gamma\mathbb{E}_{s^{\prime}\sim P(s^{\prime}|s,a)}[V^{\pi}(s^{\prime})]$ and $V^{\pi}(s)=\mathbb{E}_{a\sim\pi(a|s)}[Q(s,a)-\log\pi(a|s)]$. We see that ${\mathcal{T}}^{\pi}[Q](s,a)$ is linear in $Q$ for any $(s,a)$. Thus, the first and second terms of $J(Q,\pi)$ in (9) are linear in $Q$. The last term of (9) involves a sum of squares of linear functions of $Q$, which are convex. So, $J(Q,\pi)$ is concave in $Q$.
@@
$J(Q,\pi)$ may not necessarily be minimized at $\pi_{Q}$ such that $\pi_{Q}=\text{argmax}_{\pi}~{}~{}V^{\pi}(s)$, for all $s\in S$.
@@
    The inner minimization problem $\min_{\pi}\widehat{{\mathcal{H}}}(Q,\pi)$ has a unique optimal solution $\pi^{*}$ such that $\pi^{Q}=\text{argmin}_{\pi}V^{\pi}(s)$ for all $s\in S$ and
@@
$=\sum_{i\in[N]}\sum_{(s,a),(s^{\prime},a^{\prime})\in{\mathcal{D}}^{i}}(\overline{r}(s,a)-\overline{r}(s^{\prime},a^{\prime}))^{2}-\sum_{\begin{subarray}{c}h,k\in[N],h<k\\ \tau_{i}\in{\mathcal{D}}^{h},\tau_{j}\in{\mathcal{D}}^{k}\end{subarray}}\Big{(}R(\tau_{j})$
$-{\ln\left(\exp(R(\tau_{j}))+\exp(R(\tau_{i}))\right)}+\phi(\overline{r})\Big{)}$ (15)
We then see that the first term is a sum of squares of linear functions of $\overline{r}$, thus is strictly convex. Moreover, since $R(\tau_{i})$ is linear in $\overline{r}$ for any $\tau_{i}$ , the term $\ln(\exp(R(\tau_{i}))+\exp(R(\tau_{j})))$ has a log-sum-exp form. So this term is convex as well [3]. Putting all together we see that ${\mathcal{L}}(\overline{r})$ is strictly convex in $\overline{r}$ as desired. ∎
@@
Let $\widehat{Q}=\text{argmax}_{Q}\widehat{{\mathcal{H}}}(Q,\pi)$ and $\widehat{Q}^{C}=\text{argmax}_{Q}\widehat{{\mathcal{H}}}^{C}(Q,\pi)$, we have
@@
$\stackrel{{\scriptstyle(a)}}{{\geq}}-\beta\sum_{\begin{subarray}{c}s\sim{\mathcal{D}}\\ a\sim\mu(a|s)\end{subarray}}\widehat{Q}(s,a)+\widehat{{\mathcal{H}}}(\widehat{Q},\pi)$
@@
Appendix B Additional Details
B.1 Practical Training Objective
We discuss a practical implementation of the training objective in (4) using samples from both the expert and sub-optimal demonstration sets. We first note that the term $\mathbb{E}_{\rho_{\pi}}[{\mathcal{T}}^{\pi}[Q](s,a))]-{\mathbb{E}_{\rho_{\pi}}[\log\pi(s,a)]}$ can be written as [13]:
$\mathbb{E}_{\rho_{\pi}}[{\mathcal{T}}^{\pi}[Q](s,a))]\!-\!{\mathbb{E}_{\rho_{\pi}}[\log\pi(s,a)]}=\mathbb{E}_{(s,s^{\prime})\sim\rho^{*}}[V^{\pi}(s)-\gamma\mathbb{E}_{s^{\prime}}V^{\pi}(s^{\prime})]$
$\mathbb{E}_{\rho_{\pi}}[{\mathcal{T}}^{\pi}[Q](s,a))]-{\mathbb{E}_{\rho_{\pi}}[\log\pi(s,a)]}=\sum_{i\in[N]}w_{i}\mathbb{E}_{\rho^{i}}[V^{\pi}(s)-\gamma\mathbb{E}_{s^{\prime}}V^{\pi}(s^{\prime})]$ (18)
$-\sum_{i\in[N]}w_{i}\mathbb{E}_{\rho^{i}}[V^{Q}(s)-\gamma\mathbb{E}_{s^{\prime}}V^{Q}(s^{\prime})]-\alpha\sum_{i\in[N]}w_{i}\mathbb{E}_{\rho^{i}}[\Delta^{Q}_{\overline{r}}(s,a)]$ (19)
where $\Delta^{Q}_{\overline{r}}(s,a)=(Q(s,a)-\overline{r}(s,a))^{2}+(\mathbb{E}_{s^{\prime}}V^{Q}(s^{\prime}))^{2}+2\text{ReLU}(\overline{r}(s,a)-Q(s,a))\mathbb{E}_{s^{\prime}}V^{Q}(s^{\prime}).$
In an empirical offline implementation, to maximize $\widehat{\Gamma}(Q)$, samples $(s,a,s^{\prime})$ from demonstrations can be used to approximate the expectations over $\rho^{i}$: $\sum_{(s,a,s^{\prime})\in{\mathcal{D}}^{i}}[Q(s,a)-\gamma V^{Q}(s^{\prime})]$, $\sum_{(s,a,s^{\prime})\in{\mathcal{D}}^{i}}[V^{Q}(s)-\gamma V^{Q}(s^{\prime})]$, and $\sum_{(s,a)\in{\mathcal{D}}^{i}}[\Delta^{Q}_{\overline{r}}(s,a)]$.
We note that in continuous-action controls, the computation of $V^{Q}(s)$ involves a sum over infinitely many actions, which is impractical. In this case, we can update both $Q$ and $\pi$ in a soft actor-critic (SAC) manner. That is, for each $\pi$, we update $Q$ towards $\max_{Q}\{\widehat{{\mathcal{H}}}(Q,\pi)\}$ and for each $Q$, we update $\pi$ to bring it towards $\pi^{Q}$ by solving $\max_{\pi}\{V^{\pi}(s)\}$, for all $s$. As shown above, ^$\widehat{{\mathcal{H}}}(Q,\pi)$ in concave in $Q$ and $V^{\pi}(s)$ is convex in $\pi$, so we can expect this SAC will exhibit good behavior and stability.
B.2 Algorithm Overview
The Figure 2 provide the overview of our algorithm.
Figure 2: Overview of SPRINQL.
B.3 Environments
This section provide a detailed descrition of the enviroments used in our experiments.
B.3.1 Mujoco
MuJoCo gym environments [32] like HalfCheetah, Ant, Walker2d, Hopper, and Humanoid are integral to the field of reinforcement learning (RL), particularly in the domain of continuous control and robotics: